### v1

@@ 3.2.1 Lightweight Scheduler Design
3.2.2 Performance Functions for QoS Optimization
@@
The lightweight SCC-DSO variant replaces the full kernel regression model with the linear predictor, reducing scheduling iteration complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$. ACO pheromone updates use EWMA to avoid full matrix recalculations, further reducing overhead. The scheduler was evaluated on a 20-node Raspberry Pi Kubernetes cluster, interfaced with Prometheus for task monitoring and Grafana for latency visualization. Tests under 5G-like conditions (RTT = 40 ms, jitter = 20 ms) using Mininet-WiFi showed a 0.7% increase in mean task latency, demonstrating robustness to network variability [41].
Compared to the default Kubernetes scheduler, SCC-DSO achieved a 22.4% reduction in response time. In a city-scale smart surveillance IoT grid, it optimized video analytics workloads, improving task localization by 25% and reducing bandwidth usage by 30% on constrained nodes. The framework supports industrial integration via a RESTful API, accepting JSON-encoded job profiles and returning optimized node bindings and prefetch hints for seamless orchestration [42].
3.3 Algorithmic Complexity and Convergence
@@
4 Data Block Placement Method
4.1 Rack-Aware Data Placement Strategies
@@
4.2 Heterogeneous Sensing Data Placement
@@
Figure 3: Optimal data block placement in a heterogeneous IoT-edge cluster, balancing workloads based on node computational efficiencies (Eq. (21)). Less capable nodes receive fewer tasks, while high-performance nodes handle proportionally more.
@@
    where $m_{k}$ is the input data size (MB) for task $\text{map}_{k}$, and $T(\text{Node}_{i},\text{App}_{j},\text{map}_{k})$ is the predicted execution time (seconds) from Eq. (22). This metric quantifies node performance under ideal data locality, accounting for CPU, memory, and I/O capabilities heterogeneity.
@@
4.3 Queue-Aware Dynamic Data Placement Optimization
In perceptual cloud computing environments supporting large-scale IoT workloads, input jobs are partitioned into fixed-size blocks and distributed using multi-replica strategies to maximize fault tolerance. While this enhances data availability, it introduces task redundancy where identical blocks exist across multiple nodes, resulting in bandwidth contention, task duplication, and resource underutilization. Our proposed SCC-DSO algorithm introduces a novel hybrid scheduling approach that dynamically reorders task queues by correlating block affinity, predicted task cost, and node load, thereby minimizing redundancy and improving scheduling precision across the cluster.
@@ 4.3 Queue-Aware Dynamic Data Placement Optimization
Multidimensional performance metrics propagation delay, bandwidth, jitter, packet loss, and ACO-based cost, enable robust, real-time scheduling under strict latency and energy constraints in IoT environments for optimizing scheduling in heterogeneous sensor-cloud infrastructures. The delay function $\text{Delay}(e):E\rightarrow\mathbb{R}^{+}$ represents the expected transmission delay incurred over edge $e\in E$, where $E$ denotes the set of communication links within the cluster. This metric reflects the latency contribution of each link and is influenced by factors such as link bandwidth, queuing delay, and traffic congestion. Complementing this, the delay jitter function $\text{DelayJit}(e):E\rightarrow\mathbb{R}^{+}$ models the variability or instability of transmission delay along edge $e$, which is particularly important for real-time applications that are sensitive to timing inconsistencies [60].
@@
4.4 Adaptive Data Prefetching for Task Migration
@@
4.5 SCC-DSO Algorithm: Scheduling and Placement
@@ 5 Experimental Evaluation
SCC-DSO’s novelty lies in its integration of kernel regression for accurate execution time prediction and an adaptive task placement strategy powered by Ant Colony Optimization (ACO) with empirically tuned parameters ($\alpha=0.8$, $\beta=1.2$). This hybrid approach dynamically optimizes data locality and system resilience in heterogeneous, fluctuating environments. Experiments were conducted on a private cloud platform with 50 repetitions per configuration. Metrics were aggregated using statistical measures with 95% confidence intervals, monitored through local data placement to ensure reproducibility and transparency. Table 2 details the experimental configuration.
Table 2: Summary of Experimental Setup
@@
6 Experimental Results and Analysis
@@
Table 3: Task Completion Time Comparison (Mean $\pm$ SD) for Varying File Sizes
@@
Table 3 demonstrates that SCC-DSO consistently achieves lower task completion times than RF-FD and RSYNC across all file sizes. This improvement highlights SCC-DSO’s superior data locality and latency reduction capabilities in minimal redundancy settings.
Table 4: Task Locality Ratio (%) Across Different Cluster Sizes (Mean $\pm$ SD)
@@
Table 4 quantifies data locality improvements as cluster size increases. SCC-DSO outperforms both baselines with locality ratios exceeding 85%, confirming its ability to enhance task-data proximity and reduce cross-node data transfers in scalable environments.
@@
Table 5: Cluster Throughput (MB/s) vs. Replication Factor (Mean $\pm$ SD)
@@
Table 5 illustrates SCC-DSO’s consistently higher throughput across replication factors, indicating robustness in bandwidth utilization and fault-tolerance scenarios.
@@
Table 6: Completion Time (s) under Straggler Simulation Across Node Counts (Mean $\pm$ SD)
@@
Table 6 highlights SCC-DSO’s ability to mitigate straggler impact, resulting in substantially lower completion times, thus supporting scalability and robustness in large-scale clusters.
Table 7: Comparative Analysis of SCC-DSO and Baseline Scheduling Algorithms
@@
Table 7 summarizes the performance of SCC-DSO against RF-FD, RSYNC, and RL-Sched across key metrics. SCC-DSO achieves the highest data locality (93.1%) and lowest execution time, with excellent scalability and adaptability. Its hybrid RL+ACO design ensures efficient scheduling with moderate complexity, outperforming existing approaches in dynamic IoT-cloud environments.
@@
Figure 8 visually contrasts the evaluated frameworks, reaffirming SCC-DSO’s superior balance of performance, scalability, and efficiency. Notably, its moderate complexity makes it practical for real-world deployment in dynamic IoT-cloud infrastructures. Finally, the comprehensive experimental evaluation confirms that SCC-DSO outperforms current state-of-the-art scheduling algorithms by intelligently optimizing data locality, throughput, and execution latency, while maintaining robust scalability and adaptability for heterogeneous and large-scale cloud environments.
6.1 SCC-DSO performance under single-copy conditions
This subsection evaluates SCC-DSO in low-cost and high-overhead scenarios, focusing on a single-copy HDFS replication setting within a heterogeneous IoT cluster connected via Gigabit Ethernet. Each test was conducted 50 times, and results reflect the average metrics. SCC-DSO achieves a 13% reduction in execution time compared to RF-FD and a 7% gain over RSYNC under low network load. These improvements underscore SCC-DSO’s efficiency in environments with limited network contention [66].
SCC-DSO, by contrast, integrates a predictive performance model that accurately profiles node compute capabilities and allocates data accordingly. It dynamically adjusts scheduling by monitoring runtime queue states and proactively triggers data prefetching for tasks anticipated to execute on non-local nodes. This preemptive strategy reduces idle time and mitigates bandwidth contention. The resultant alignment between data locality and node performance ensures reduced execution time, minimized data migration, and improved throughput under single-copy storage constraints. These findings validate SCC-DSO’s efficacy in optimizing data placement and scheduling in bandwidth-rich, compute-diverse IoT environments [66].
@@ 6.1 SCC-DSO performance under single-copy conditions
6.2 SCC-DSO performance in multi-copy conditions
@@
This preemptive data staging effectively decouples computation from I/O latency, sustaining pipeline throughput and alleviating backpressure on high-performance nodes. Consequently, SCC-DSO incorporates a decentralized, redundant, and aware task scheduler that leverages temporal locality and multi-copy data placement heuristics to maximize intra-node data reuse. We maintain consistent input availability near task invocation windows, which enables the scheduler to achieve high core utilization and effectively mitigate bandwidth-induced stalls. Collectively, these advances establish SCC-DSO as a robust solution for throughput optimization in distributed IoT clusters characterized by network sensitivity and data redundancy [67].
@@
Figure 10 illustrates the variation in data execution rate (%) concerning increasing data block sizes under different synchronization thresholds ($\theta=\{0.2,0.4,0.6,0.8\}$). SCC-DSO demonstrates consistently higher data execution rates across all scenarios, suggesting an improved ability to maintain execution fidelity even as synchronization pressure increases. Notably, at lower $\theta$ values, the gap in performance between SCC-DSO and the baseline methods (RF-FD and RSYNC) is pronounced, reflecting SCC-DSO’s capability to optimize partial state synchronization and exploit semantic-aware delta selection. As $\theta$ approaches 0.8, the execution rates of all methods converge; however, SCC-DSO reaches near-optimal performance ($\geq 99\%$), indicating its resilience to state divergence and minimal rollback overhead. This superior performance can be attributed to SCC-DSO’s dynamic state consistency control and its likely use of asynchronous conflict resolution, enabling high execution throughput without sacrificing consistency guarantees. These results validate that SCC-DSO not only reduces latency (as shown in Figure 9) but also sustains high execution rates, thereby enhancing the end-to-end efficiency of distributed systems operating under large-scale and weakly consistent constraints.
7 Limitations and Future Work
Despite SCC-DSO’s demonstrated efficiency, several open challenges remain. First, its hybrid use of ant colony optimization and kernel regression introduces non-trivial computational complexity, limiting deployment on ultra-low-power edge devices. Future work will explore neuromorphic and inspired surrogates, low-rank approximations, and hardware-aware pruning to enable sub-second inference within microcontroller and class IoT nodes [68]. Second, current validation is confined to a homogeneous 50-node cluster under deterministic 1 Gbps Ethernet; this leaves open questions regarding SCC-DSO’s scalability and fault-tolerance under geo-distributed, bandwidth-fluctuating, and latency-sensitive edge-cloud topologies. Moreover, the absence of cross-layer adaptivity restricts responsiveness to volatile wireless substrates such as 5G NR, LPWAN, or mobile fog fabrics. To address this, future research will embed self-adaptive policy networks using meta-reinforcement learning, integrate decentralized orchestration layers (e.g., KubeEdge, WebAssembly), and couple real-time telemetry feedback with task migration strategies. Experimental deployment over federated testbeds (e.g., Fed4FIRE+, 5TONIC) will serve to validate robustness and energy-performance tradeoffs in next-generation IoT-cloud continuums [68].
8 Conclusion
@@
This paper investigates performance degradation of jobs during execution in heterogeneous IoT clusters and proposes SCC-DSO, a perceptual cloud-based data scheduling optimization framework. However, SCC-DSO adaptively allocates heterogeneous-sized data blocks to working nodes by real-time profiling their computational capabilities, leveraging predictive performance modeling to guide scheduling decisions. Firstly, the framework introduces dynamic data migration to maximize localized execution, minimizing data transfer overhead and network latency. Consequently, central to SCC-DSO is a data scheduling queue optimization algorithm, which utilizes initial scheduling lists to construct an optimized pre-allocation queue, enabling parallel, contention-free local scheduling across distributed nodes. Secondly, a novel prefetching strategy based on minimum task queue completion time is integrated, effectively overlapping computation and communication to reduce idle times associated with non-local data dependencies. The algorithm also explicitly accounts for data reliability by considering replication and fault tolerance in scheduling decisions. Thirdly, comprehensive experiments demonstrate that SCC-DSO achieves up to 19.8% and 7.6% reductions in execution time compared to RF-FD and RSYNC benchmarks, respectively, significantly improving data locality and throughput in multi-copy scenarios. Finally, these results validate SCC-DSO’s efficacy in optimizing data scheduling to enhance resource utilization and job performance in complex IoT-cloud environments characterized by heterogeneity and dynamic workloads.

### v2

@@ III-B1 Lightweight Scheduler Design
III-B2 Performance Functions for QoS Optimization
@@
The lightweight SCC-DSO variant replaces the full kernel regression model with the linear predictor, reducing scheduling iteration complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$. ACO pheromone updates use EWMA to avoid full matrix recalculations, further reducing overhead. The scheduler was evaluated on a 20-node Raspberry Pi Kubernetes cluster, interfaced with Prometheus for task monitoring and Grafana for latency visualization. Tests under 5G-like conditions (RTT = 40 ms, jitter = 20 ms) using Mininet-WiFi showed a 0.7% increase in mean task latency, demonstrating robustness to network variability [41]. SCC-DSO achieved a 22.4% mitigation time compared with the default Kubernetes scheduler optimization. In a city-scale smart surveillance IoT grid, it optimized video analytics workloads, improving task localization by 25% and reducing bandwidth usage by 30% on constrained nodes. The framework supports industrial integration via a RESTful API, accepting JSON-encoded job profiles and returning optimized node bindings and prefetch hints for seamless orchestration [42].
III-C Algorithmic Complexity and Convergence
@@
IV Data Block Placement Method
IV-A Rack-Aware Data Placement Strategies
@@
IV-B Heterogeneous Sensing Data Placement
@@
    where $m_{k}$ is the input data size (MB) for task $\text{map}_{k}$, and $T(\text{Node}_{i},\text{App}_{j},\text{map}_{k})$ is the predicted execution time (seconds) from Eq. (19). This metric quantifies node performance under ideal data locality, taking into account the heterogeneity of CPU, memory, and I/O capabilities.
@@
Figure 4: Optimal data block placement in a heterogeneous IoT-edge cluster, balancing workloads based on node computational efficiencies (Eq. (18)). Grid-based block partitioning strategy for parallel spatial data processing.
Figure 4 illustrates that a grid-based block partitioning strategy is widely adopted in parallel and distributed computing for spatial data analysis, matrix computation, and domain decomposition techniques. The grid is divided into nine sub-blocks $B_{1}$ through $B_{9}$, representing distinct portions of the computational domain. The color scheme implies functional segregation: the gray blocks ($B_{1}$$B_{4}$) denote initial or pre-processed data regions, the blue blocks ($B_{5}$$B_{6}$) correspond to core computation zones, while the green blocks ($B_{7}$$B_{9}$) represent final or post-processing regions. This hierarchical decomposition facilitates enhanced computational efficiency, optimized memory locality, and balanced workload distribution—critical components in high-performance data processing. Moreover, the structured layout supports region-wise parallel execution and reduces inter-process communication overhead, offering scalable performance for large-scale scientific simulations and AI-driven spatial modeling applications.
IV-C Queue-Aware Dynamic Data Placement Optimization
In perceptual cloud computing environments supporting large-scale IoT workloads, input jobs are partitioned into fixed-size blocks and distributed using multi-replica strategies to maximize fault tolerance. While this enhances data availability, it introduces task redundancy where identical blocks exist across multiple nodes, resulting in bandwidth contention, task duplication, and resource underutilization. Our proposed SCC-DSO algorithm introduced a novel hybrid scheduling approach that dynamically reorders task queues by correlating block affinity, predicted task cost, and node load, thereby minimizing redundancy and improving scheduling precision across the cluster.
Conversely, multidimensional performance metrics propagation delay, bandwidth, jitter, packet loss, and ACO-based cost, enable robust, real-time scheduling under strict latency and energy constraints in IoT environments for optimizing scheduling in heterogeneous sensor-cloud infrastructures. The delay function $\text{Delay}(e):E\rightarrow\mathbb{R}^{+}$ represents the expected transmission delay incurred over edge $e\in E$, where $E$ denotes the set of communication links within the cluster. This metric reflects the latency contribution of each link and is influenced by factors such as link bandwidth, queuing delay, and traffic congestion. Complementing this, the delay jitter function $\text{DelayJit}(e):E\rightarrow\mathbb{R}^{+}$ models the variability or instability of transmission delay along edge $e$, which is particularly important for real-time applications that are sensitive to timing inconsistencies [46].
@@ IV-C Queue-Aware Dynamic Data Placement Optimization
IV-D Adaptive Data Prefetching for Task Migration
@@
IV-E SCC-DSO Algorithm: Scheduling and Placement
@@ V Experimental Evaluation
TABLE II: Summary of Experimental Setup
@@
SCC-DSO’s novelty lies in its integration of kernel regression for accurate execution time prediction and an adaptive task placement strategy powered by Ant Colony Optimization (ACO) with empirically tuned parameters ($\alpha=0.8$, $\beta=1.2$). This hybrid approach dynamically optimizes data locality and system resilience in heterogeneous, fluctuating environments. Experiments were conducted on a private cloud platform with 50 repetitions per configuration. Metrics were aggregated using statistical measures with 95% confidence intervals, monitored through local data placement to ensure reproducibility and transparency. Table II details the experimental configuration.
VI Experimental Results and Analysis
@@
TABLE III: Task Completion Time Comparison (Mean $\pm$ SD) for Varying File Sizes
@@
Table III demonstrates that SCC-DSO consistently achieves lower task completion times than RF-FD and RSYNC across all file sizes. This improvement highlights SCC-DSO’s superior data locality and latency reduction capabilities in minimal redundancy settings.
TABLE IV: Task Locality Ratio (%) Across Different Cluster Sizes (Mean $\pm$ SD)
@@
Table IV quantifies data locality improvements as cluster size increases. SCC-DSO outperforms both baselines with locality ratios exceeding 85%, confirming its ability to enhance task-data proximity and reduce cross-node data transfers in scalable environments.
@@
TABLE V: Cluster Throughput (MB/s) vs. Replication Factor (Mean $\pm$ SD)
@@
Table V illustrates SCC-DSO’s consistently higher throughput across replication factors, indicating robustness in bandwidth utilization and fault-tolerance scenarios.
@@
TABLE VI: Completion Time (s) under Straggler Simulation Across Node Counts (Mean $\pm$ SD)
@@
Table VI highlights SCC-DSO’s ability to mitigate straggler impact, resulting in substantially lower completion times, thus supporting scalability and robustness in large-scale clusters.
TABLE VII: Comparative Analysis of SCC-DSO and Baseline Scheduling Algorithms
@@
Table VII summarizes the performance of SCC-DSO against RF-FD, RSYNC, and RL-Sched across key metrics. SCC-DSO achieves the highest data locality (93.1%) and lowest execution time, with excellent scalability and adaptability. Its hybrid RL+ACO design ensures efficient scheduling with moderate complexity, outperforming existing approaches in dynamic IoT-cloud environments.
@@
VI-A SCC-DSO performance under single-copy conditions
This subsection evaluates SCC-DSO in low-cost and high-overhead scenarios, focusing on a single-copy HDFS replication setting within a heterogeneous IoT cluster connected via Gigabit Ethernet. Each test was conducted 50 times, and results reflect the average metrics. SCC-DSO achieves a 13% reduction in execution time compared to RF-FD and a 7% gain over RSYNC under low network load. These improvements underscore SCC-DSO’s efficiency in environments with limited network contention [66]. SCC-DSO, by contrast, integrates a predictive performance model that accurately profiles node compute capabilities and allocates data accordingly. It dynamically adjusts scheduling by monitoring runtime queue states and proactively triggers data prefetching for tasks anticipated to execute on non-local nodes. This preemptive strategy reduces idle time and mitigates bandwidth contention. The resultant alignment between data locality and node performance ensures reduced execution time, minimized data migration, and improved throughput under single-copy storage constraints. These findings validate SCC-DSO’s efficacy in optimizing data placement and scheduling in bandwidth-rich, compute-diverse IoT environments [66].
@@ VI-A SCC-DSO performance under single-copy conditions
VI-B SCC-DSO performance in multi-copy conditions
@@
This preemptive data staging effectively decouples computation from I/O latency, sustaining pipeline throughput and alleviating backpressure on high-performance nodes. Consequently, SCC-DSO incorporates a decentralized, redundancy-aware task scheduler that leverages temporal locality and multi-copy data placement heuristics to maximize intra-node data reuse. We maintain consistent input availability near task invocation windows, which enables the scheduler to achieve high core utilization and effectively mitigate bandwidth-induced stalls. Collectively, these advances establish SCC-DSO as a robust solution for throughput optimization in distributed IoT clusters characterized by network sensitivity and data redundancy [67].
@@
Figure 11 illustrates the variation in data execution rate (%) concerning increasing data block sizes under different synchronization thresholds ($\theta=\{0.2,0.4,0.6,0.8\}$). SCC-DSO demonstrates consistently higher data execution rates across all scenarios, suggesting an improved ability to maintain execution fidelity even as synchronization pressure increases. Notably, at lower $\theta$ values, the gap in performance between SCC-DSO and the baseline methods (RF-FD and RSYNC) is pronounced, reflecting SCC-DSO’s capability to optimize partial state synchronization and exploit semantic-aware delta selection. As $\theta$ approaches 0.8, the execution rates of all methods converge; however, SCC-DSO reaches near-optimal performance ($\geq 99\%$), indicating its resilience to state divergence and minimal rollback overhead. This superior performance can be attributed to SCC-DSO’s dynamic state consistency control and likely use of asynchronous conflict resolution, enabling high execution throughput without sacrificing consistency guarantees.
VII Limitations and Future Work
Privacy concerns arise as SCC-DSO processes sensitive IoT data such as location and health metrics. To mitigate these risks, we propose incorporating differential privacy into the Gaussian Process Regression (GPR) model by injecting Gaussian noise, $\varepsilon\sim\mathcal{N}(0,0.01)$, into execution time predictions. Formally, the GPR prediction model can be expressed as
$y_{i}=f(x_{i})+\varepsilon,\quad f\sim\mathcal{GP}(0,k(x,x^{\prime})),$
where $f$ is a Gaussian process with kernel $k$, and $\varepsilon$ is Gaussian noise. This approach balances privacy preservation with scheduling accuracy. Additionally, data block placement strategies (Section IV) utilize encrypted HDFS storage and secure MQTT brokers, complying with GDPR and IoT security standards to prevent data leakage in distributed environments [68]. Regarding sustainability, SCC-DSO’s lightweight RL-ACO variant, with computational complexity $\mathcal{O}(n\log n)$, achieves approximately 15% energy savings compared to DQN-RL baselines on a 50-node cluster, attributed to reduced iteration counts (12 vs. 20). Future work includes precise carbon footprint quantification via GreenCloud simulations, aiming for carbon-neutral scheduling by leveraging renewable-powered nodes. Moreover, exploring neuromorphic surrogate models promises accelerated convergence and further energy efficiency improvements. These directions underscore SCC-DSO’s potential for sustainable, privacy-aware IoT-cloud deployments.
VIII Conclusion
@@
This paper presents SCC-DSO, a novel perceptual cloud-based data scheduling optimization framework designed to mitigate performance degradation in heterogeneous IoT clusters. SCC-DSO adaptively partitions and allocates heterogeneous-sized data blocks to compute nodes by real-time profiling of their computational capacities, guided by predictive performance models. Key innovations include a dynamic data migration mechanism that maximizes localized execution, reducing network latency and data transfer overhead. Data scheduling queue optimization algorithms rely on initial schedules to construct content-free queues that enable efficient parallel local scheduling across distributed nodes. Moreover, by minimizing the completion time of task queues, a novel way of prefetching tasks effectively overlaps computation and communication, reducing idle periods associated with data dependencies. SCC-DSO also integrates data reliability considerations by incorporating replication and fault tolerance into scheduling decisions. Extensive experiments demonstrate up to 19.8% and 7.6% reductions in execution time compared to RF-FD and RSYNC benchmarks, respectively, while significantly enhancing data locality and throughput under multi-copy scenarios. These results validate SCC-DSO’s capability to optimize resource utilization and improve job performance in complex, dynamic, and heterogeneous IoT-cloud environments, advancing the state of scalable, adaptive data scheduling.
Appendix A Notation Summary
Table IX provides clear and consistent definitions of symbols and variables.
TABLE IX: Notation Summary for SCC-DSO Framework
Symbol Definition
$V$ Set of computing nodes in the IoT-edge cluster
$E$ Set of communication links between nodes
$J,T$ Set of tasks
$x_{ij}$ Binary variable indicating assignment of task $j$ to node $i$
$T_{ij}$ Predicted execution time of task $j$ on node $i$
$d_{j}$ Data size required by task $j$
$C_{i}(t)$ Dynamic computational capacity of node $i$ at time $t$
$\tau_{ij}$ Pheromone trail value for task $j$ on node $i$
$\eta_{ij}$ Heuristic desirability for assigning task $j$ to node $i$
$\alpha$ Pheromone influence parameter
$\beta$ Heuristic influence parameter
$\rho$ Pheromone evaporation rate
$Q$ Pheromone deposit constant
$K$ Number of ants in ACO
II Maximum iterations in ACO
$w_{1},w_{2},w_{3}$ Weight factors for delay, energy, and packet loss
$\phi$ Prefetch threshold
$\gamma$ Learning rate for kernel regression model
$\sigma$ RBF kernel bandwidth
$\theta$ Task migration limit per iteration
$\epsilon$ Convergence tolerance for ACO
$\delta$ Minimum pheromone trail
$B$ Data block size or total number of blocks (context-dependent)
$P(Node_{i},App_{j},map_{k})$ Computational efficiency of node $i$
PLFPLF Prefetch Load Factor between target and source node
$RQ_{i}$ Resource Quotient for node $i$
$Locality_{j,i}$ Local data presence indicator for task $j$ on node $i$
$Eff_{j,i}$ Efficiency metric combining locality and execution time
$BW_{x}$ Bandwidth capacity of node or edge $x$
$LossPkt(v)$ Packet loss probability at node $v$